### v5

# Knowledge Transfer from High-Resource to Low-Resource Programming Languages for Code LLMs

Abstract.
Over the past few years, Large Language Models of Code (Code LLMs) have started to have a significant impact on programming practice. Code LLMs are also emerging as building blocks for research in programming languages and software engineering. However, the quality of code produced by a Code LLM varies significantly by programming language. Code LLMs produce impressive results on programming languages that are well represented in their training data (e.g., Java, Python, or JavaScript), but struggle with low-resource languages that have limited training data available. Low resource languages include OCaml, Racket, and several others.
This paper presents an effective approach for boosting the performance of Code LLMs on low-resource languages using semi-synthetic data. Our approach generates high-quality datasets for low-resource languages, which can then be used to fine-tune any pretrained Code LLM. Our approach, called MultiPL-T, translates training data from high-resource languages into training data for low-resource languages in the following way. 1) We use a Code LLM to synthesize tests for commented code from a high-resource language, filtering out faulty tests and code with low test coverage. 2) We use a Code LLM to translate Python code to a target low-resource language, and use tests to validate the translation. We apply this approach to generate tens of thousands of new, validated training items for Julia, Lua, OCaml, R, and Racket. Furthermore, we use an open model (StarCoderBase) with open training data (The Stack), which allows us to decontaminate benchmarks, train models without violating licenses, and run experiments that could not otherwise be done.
With MultiPL-T generated data, we present fine-tuned versions of StarCoderBase and Code Llama for Julia, Lua, OCaml, R, and Racket. On established benchmarks (MultiPL-E), these models outperform other open Code LLMs. The MultiPL-T approach is easy to apply to new languages, and is significantly more efficient and effective than alternatives such as training longer.
copyright: none
1. Introduction
Large Language Models of Code (Code LLMs) are starting to have a significant impact on both professional programmers and research in programming languages and software engineering. GitHub Copilot is just one of several popular tools powered by Code LLMs (Copilot, 2023; CodeWhisperer, 2023; TabNine, 2023). Moreover, Code LLMs are also emerging as a building block for research (Schäfer et al., 2023; Bareiß et al., 2022; Murali et al., 2023; Nam et al., 2023; Ross et al., 2023; Lemieux et al., 2023; Xia et al., 2023; Joshi et al., 2023; Phung et al., 2023; First et al., 2023; Chen et al., 2023). However, the quality of code produced by a Code LLM varies significantly by programming language. Models are most impressive at high-resource programming languages such as Python, JavaScript, and Java, but struggle to produce code in low-resource languages, such as Racket and OCaml (Cassano et al., 2023; Athiwaratkun et al., 2022; Zheng et al., 2023). This puts programmers who rely on these languages at a disadvantage, since they do not receive the same benefits that Code LLMs can deliver for high-resource languages (Ziegler et al., 2022; Murali et al., 2023).
The key issue is that the performance of Code LLMs depends on the amount of language data available for training. For example, The Stack, which is the training set for several contemporary Code LLMs, has 64GB of Python, but only around 1GB of OCaml and 0.5GB of Scheme/Racket (Kocetkov et al., 2023). [^1] This is the volume of data that remains after files are deduplication for training (Li et al., 2023). As figure 1 shows, the performance of StarCoderBase, an open Code LLM trained on The Stack, generally increases as the proportion of training data for the language increases.
Figure 1. The performance of StarCoderBase-15B on several languages from the MultiPL-E benchmark for Code LLMs, plotted against their proportion of the model’s training data. Using MultiPL-T, this paper significantly improves how StarCoderBase-15B performs on several low-resource languages, as shown by the arrows. The bottom of each arrow indicates how the base model performs, and the arrowheads indicate performance after fine-tuning with MultiPL-T.
Our goal in this paper is to investigate methods to further train, or fine-tune, pretrained Code LLMs to improve their performance on low-resource languages. The obvious approach is try to find more data, but for a low-resource language, more data is hard to find by definition. For example, The Stack already includes all permissively licensed code for 358 programming languages from GitHub as of 2022, and GitHub is by far the largest repository of open source code (Borges et al., 2016). [^2] The Stack deliberately excludes copyleft licenses and unlicensed code. An alternative is to train longer (i.e, for more epochs) on existing data. However, in this paper we show that training longer on several low-resource languages is not only inefficient, but can actually hurt performance (§ 3.1). Another alternative is to train models on synthetic data. Self-instruction (Wang et al., 2023) and instruction evolution (Luo et al., 2023) fine-tune LLMs on data generated by another LLM. These approaches have been effective at fine-tuning models on high-resource programming languages. However, we show that self-instruction does not work for low-resource languages, because for low-resource languages, the LLM-generated programs are of poor quality (§ 3.2). Here, ‘quality’ refers to both the correctness of the code (i.e., its ability to function as intended) and its stylistic integrity (i.e., adherence to best practices and readability).
In this paper, we present a new and effective approach for fine-tuning Code LLMs for low-resource programming languages that is based on generating semi-synthetic training data. Our approach relies on several key ingredients. 1) The large volume of training data for high-resource programming languages includes a lot of well-documented code; 2) Code LLMs are effective and efficient unit test generators, and we can check that generated tests are valid (Schäfer et al., 2023); 3) We can mechanically translate many unit tests to a low-resource language with a simple compiler (Roziere et al., 2021; Athiwaratkun et al., 2022; Cassano et al., 2023); 4) Code LLMs can translate code from one language to another, and we can test these translations with the aforementioned tests, retry until tests pass, and engineer a prompt to increase the likelihood of a successful translation. Putting these four ideas together, we develop a pipeline for transferring training data across multiple programming languages that we call MultiPL-T.
Using training data generated by MultiPL-T, we present fine-tuned Code LLMs that achieve state-of-the-art performance on five low-resource languages: Racket, OCaml, Lua, R, and Julia. We focus primarily on fine-tuning the StarCoder family of Code LLMs (Li et al., 2023). There are StarCoder models available at a variety of sizes, including a 1B parameter model that is lightweight enough to run on CPUs, and a more capable 15B parameter model, which we use as the test generator and language translator for MultiPL-T. The StarCoder models also have open training data, which allows us to compare MultiPL-T to a baseline of training longer on existing data for low-resource languages. Finally, we also present fine-tuned versions of the Code Llama 34B (Rozière et al., 2023) model, which was recently released.
Contributions
To summarize, we make the following contributions:
  1. (1)
    MultiPL-T, an effective approach for generating semi-synthetic data for low-resource programming languages using test-validated translation of high-quality code in high-resource languages.
  2. (2)
    Efficient fine-tuning datasets for Julia, Lua, OCaml, R, and Racket, comprising tens of thousands of documented and tested functions.
  3. (3)
    A dataset of 133,168 Python functions extracted from the Stack, where every function has natural language documentation and a validated set of generated tests with high coverage. This dataset could be used to generate fine-tuning sets for other programming languages.
  4. (4)
    Fine-tuned versions of StarCoderBase and Code Llama for Julia, Lua, OCaml, R, and Racket in three sizes: 1B, 15B, and 34B parameters. The 1B models can be run relatively well on CPUs, whereas the larger models require GPUs to produce results with reasonable latency.
  5. (5)
    A thorough evaluation that includes a) a comparison of MultiPL-T to the baseline of training further on existing data, b) an evaluation of the fine-tuning efficiency with MultiPL-T, c) results on prior multi-language benchmarks (Cassano et al., 2023), d) a new multi-language benchmark designed to exercise in-context learning, e) an evaluation of how generated code adheres to common Racket programming style.
The MultiPL-T datasets, models, and the code to reproduce our work are available at:
https://huggingface.co/datasets/nuprl/MultiPL-T
2. Background
In this section, we give a high-level overview of how Code LLMs are trained and evaluated. We use StarCoder as the example, since it is the model that we use for most of MultiPL-T.
Training Large Language Models of Code
A large language model (LLM) is a neural network trained on hundreds of gigabytes or even terabytes of data. Code LLMs are trained on source code (and often natural language documents too), which allows them to generate code from comments, comments from code, more code from code, and so on. LLM training takes significant resources: StarCoderBase was trained on approximately 800GB of code, which took three weeks on a cluster of 512 NVIDIA A100 GPUs.
The only way to build a training set of this scale is to scrape public repositories of code. There are a handful of public training sets that are based on GitHub (Xu et al., 2022; Felipe Hoffa, 2016), but the largest and most recent is The Stack (Kocetkov et al., 2023). The Stack v1.2 has 3TB of permissively licensed source code for 358 programming languages. It was constructed in 2022, and has since been used to train several Code LLMs (Allal et al., 2023; Replit, 2023; Nijkamp et al., 2023), including StarCoderBase. Specifically, StarCoderBase was trained on a subset of The Stack consisting of 86 programming languages.
def vowels_count(s):
    """Write a function vowels_count which takes a string representing a word as
    input and returns the number of vowels in the string.  Vowels in this case are
    a’, e’, i’, o’, u’. Here, y is also a vowel, but only when it is at the
    end of the given word.
(a) Python prompt.
(* Write a function vowels_count which takes a string representing a word as
   input and returns the number of vowels in the string.  Vowels in this case are
   a’, e’, i’, o’, u’. Here, y is also a vowel, but only when it is at the
   end of the given word.  *)
let vowels_count (s : str) : int =
(b) OCaml prompt.
Figure 2. An example prompt from a HumanEval problem and its translation to OCaml, with our extension to MultiPL-E. Not shown are doctests and hidden test cases, which are also translated to OCaml. This particular problem is hard for many LLMs because it alters the strong prior on what vowels are, by saying that y is a vowel when it is the last letter in a word.
The StarCoder model family
StarCoder is a family of models that are available at several sizes (Li et al., 2023). The largest and most capable model is the family is called StarCoderBase: it has 15B parameters, which requires a capable GPU to run inference with reasonable latency. There are smaller versions of StarCoderBase that were trained on exactly the same data. To make use of limited GPU resources, we use the smallest model, StarCoderBase-1B, for most experiments in this paper. However, we also show that our results generalize to StarCoderBase-15B. There is also a model in the StarCoder family that is just named StarCoder: it is StarCoderBase-15B specialized to excel at Python. [^3] StarCoder fine-tunes StarCoderBase-15B on two more epochs of Python data from The Stack. This paper uses StarCoderBase-15B for translations to low-resource languages, and StarCoder-15B for Python test generation.
The Code Llama model family
The Code Llama (Rozière et al., 2023) family of models were recently released and perform better than the StarCoder models on common benchmarks. While the authors state that the training data comes from publicly accessible datasets, they do not disclose the specific datasets used, preventing us from conducting the exhaustive evaluation on Code Llama that we do with StarCoder and its training data. Moreover, the Llama license forbids using model outputs to train non-Llama models, which is why MultiPL-T uses StarCoder for data generation. However, we train and evaluate the largest Code Llama on data constructed with StarCoder models.
Fine-tuning
After training, a model can be further trained, or fine-tuned, with significantly fewer resources. For example, there are several fine-tuned versions of StarCoderBase that were trained with a few days of GPU time on a modest amount of data (e.g., (Luo et al., 2023; Muennighoff et al., 2023)). Most fine-tuned versions of StarCoderBase are designed to make the model even better at high-resource languages, such as Python. In contrast, this paper presents fine-tuned versions of StarCoderBase that are significantly better at several low-resource languages.
Benchmarking Code LLMs
Most Code LLM benchmarks, including those we use in this paper, follow the format introduced by the Codex “HumanEval” benchmark (Chen et al., 2021). Every benchmark problem has two parts: 1) a prompt for the LLM that has a function signature and a comment, and 2) a suite of test cases that are not given to the LLM. Thus each problem is run in two steps: 1) the model generates a function from the prompt, and 2) the generated function is then tested with the hidden tests, and all tests must pass for the generated code to be considered correct.
[…]
Code LLMs appear to produce higher quality code when their output is sampled (Chen et al., 2021). Since sampling introduces non-determinism, we must evaluate their output by generating several samples from the same prompt. The most widely used metric for Code LLM performance is pass@k, which is the likelihood that the LLM produces a correct program at least once from $k$ attempts. Pass@k must be estimated from $n>>k$ samples. When $k=1$, pass@1 is the same as the pass rate ($k/n$). We use pass@1 as the metric for all our benchmarking experiments, which is common practice. Intuitively, pass@1 measures the ability of the Code LLM to generate a correct solution in a single attempt.
3. Alternatives to MultiPL-T
Before we present the MultiPL-T approach, we consider two simpler alternatives.
(a) Fine-tuning on the complete language
specific subsets of The Stack.
(b) Fine-tuning on subsets that are approximately the same size as the MultiPL-T fine-tuning datasets.
Figure 3. We fine-tune StarCoderBase-1B on several epochs of language-specific data of The Stack and measure performance with MultiPL-E. In figure 2(a), we train on all data from The Stack for each language. These datasets vary in size (the labels measure their size in tokens). In figure 2(b) we sample each dataset to be approximately the same size as the MultiPL-T datasets. Both approaches that use data from The Stack barely improve performance, and can even hurt performance. In contrast, fine-tuning on MultiPL-T (dashed lines) shows significant improvement.
3.1. Further Training on Natural Data
The simplest way to boost the performance of a Code LLM on a programming language is to train it further on natural data, which is code written by human programmers rather than code generated by other means (e.g., an LLM). This was the approach taken to create StarCoder from StarCoderBase. The latter is the base model, and the former is fine-tuned on roughly two additional epochs [^4] An epoch in machine learning refers to one complete pass through the entire training dataset. of the Python subset of The Stack. Although this approach is effective for high-resource languages, we now show that it does not work for several low-resource languages (Figure 3).
In figure 2(a), we fine-tune three versions of StarCoderBase-1B on three more epochs of Lua, OCaml, and Racket each. This data is from The Stack. However, The Racket subset of The Stack is poor quality, so we use the Scheme subset instead. [^5] The Racket subset accidentally omits the .rkt file extension and largely contains Racket documentation (in Scribble). Since Racket is descendant from Scheme, the Scheme subset is a more reasonable fine-tuning set. The Stack has an order of magnitude more Lua than OCaml and Racket. Moreover, even the Racket and OCaml data in The Stack is significantly larger than the fine-tuning datasets we will develop with MultiPL-T. Therefore, these experiments are not directly comparable to each other, since they train on wildly varying amounts of data. Nevertheless, we get poor results for all: the performance of these fine-tuned models barely increases for Racket and OCaml and even decreases for Lua.
In figure 2(b), we do another experiment with The Stack that lends itself to a direct comparison with MultiPL-T. We randomly sample data from The Stack to get approximately the same volume of data that we generate with MultiPL-T. Thus, fine-tuning on these datasets will use similar computing resources as fine-tuning a model with MultiPL-T data. We use this data to fine-tune three versions on StarCoderBase-1B for six epochs, and evaluate the models at each epoch. We still get poor results with The Stack: Lua and OCaml performance barely increases whereas Racket performance decreases. In contrast, fine-tuning with MultiPL-T (dashed lines) will show significant gains.
;; sum-digits-of-number: Number -> Number
;; Sums the digits of a number
(define (sum-digits-of-number num)
    (foldr + 0 (map string->number (string-split (number->string num) ""))))
(a) Splitting on the empty string produces empty strings in the list, which cannot be converted to a number.
;; add-rev-nums: [List-of Numbers] -> Number
;; Adds the reverse of all the numbers in a list
(define (add-rev-nums lon)
    (foldr + 0 (map (lambda (n)
                        (string->number (list->string
                            (reverse (string->list (number->string n))))))
                    lon)))
(b) Working function.
;; is-palindrome-sum?: Number -> Boolean
;; Checks if a number is a palindrome sum of its digits
(define (is-palindrome-sum? num)
    (is-palindrome? (sum-remainders
      (map char->integer
        (string->list (number->string num))) 10)))
(c) char->integer returns the ASCII code of a character, and does not parse it as a digit.
;; sum-palindromes: [List-of Numbers] -> Number
;; Sums the palindrome numbers in the list
(define (sum-palindromes lon)
    (foldr + 0 (filter is-palindrome? lon)))
(d) Uses a function from the prompt, despite separators.
;; is-prime?: Number -> Boolean
;; Checks if a number is prime
(define (is-prime? num)
    (cond
        ((= num 2) #t)
        ((= num 1) #f)
        (else (zero? (remainder (add-odds (repeat "2" (/ num 2))) num)))))
(e) Several errors, e.g., repeat is hallucinated and uses a function from prompt.
Figure 4. Faulty Racket code generated by StarCoderBase-15B when seeded with five hand-written examples.
3.2. Self-Instruction for Low-Resource Programming Languages
An alternative to fine-tuning on natural data is to fine-tune on LLM-generated data (Wang et al., 2023; Luo et al., 2023). The usual approach is to seed a model with hand-selected programs, prompt the model to generate more examples, and iterate until a large enough dataset has been collected. This type of approach has been used successfully to generate training data for Code LLMs in high-resource languages (Luo et al., 2023). However, we argue that self-instruction and its variations are unlikely to work for low-resource languages because Code LLMs often generate low-quality code in low-resource languages.
To illustrate, we use StarCoderBase-15B to generate functions in Racket, mimicking the first step of self instruction. We prompt the model with five hand-written examples (Appendix A) and have it generate five more functions, shown in Figure 4. Unfortunately, we find that three of the five model-generated programs have bugs, which we explain in the captions. This is a much higher error rate than what is evident from self-instruct datasets for high-resource languages (Chaudhary, 2023; Muennighoff et al., 2023).
Self-instruction and training further on existing public data do not help Code LLMs perform better on low-resource programming languages. A final alternative is to train on proprietary data, which does not lend itself to the kind of detailed evaluation that we do (§§ 3.1 and 5). Thus we now turn to the MultiPL-T approach.
Figure 5. The MultiPL-T pipeline for generating semi-synthetic data. Starting with a tested Python function \⃝raisebox{-0.9pt}{1}, we translate the header and test cases into each target language \⃝raisebox{-0.9pt}{2}. We prompt the model with the translated header and comment to generate 50-100 candidate translations \⃝raisebox{-0.9pt}{3} (varies by language). We append the compiled test cases to the candidate translations \⃝raisebox{-0.9pt}{4} and evaluate with the target language runtime. We deduplicate all programs that pass all test cases \⃝raisebox{-0.9pt}{5} to build the fine-tuning dataset \⃝raisebox{-0.9pt}{6}.
4. Our Approach
We now present the MultiPL-T approach to generating high-quality, semi-synthetic data for low-resource languages. Figure 5 depicts the MultiPL-T system, which has several stages. 1) Given a training dataset (The Stack), we filter data from a high-resource language (Python) to select code that is amenable to automatic test generation and translation. The Stack has 60GB of Python, and translation and test generation are expensive, so we filter quite aggressively. We only select individual Python functions that have docstrings and pass a heuristic type-checker (§ 4.1). 2) Given the filtered dataset, we use a Code LLM (StarCoder-15B) to generate test suites for each function. We validate the generated tests for correctness and code coverage, and find that the Code LLM can be used as a capable test generator for our purposes (§ 4.2). 3) We translate each Python function to a target language $L$, by prompting the Code LLM to translate code. This translation may go wrong, especially because the Code LLM performs poorly on the low-resource target language. 4) We filter the $L$ functions (from Step 1) to only select those that pass test cases. To do so, we compile the Python test cases (from Step 2) to the language $L$, using the Python-to-$L$ test case compiler from MultiPL-E. The test case compiler is a traditional compiler that does not suffer from LLM hallucinations: if it cannot compile a test case, it signals an error, and we discard the training item if too many test cases fail to compile (§ 4.3). The final result is thus a dataset of novel training items for the language $L$, which may be used to fine-tune any LLM. In § 5, we discuss how we use this data to fine-tune and evaluate several models for five different low-resource languages. The rest of this section describes the above steps in depth.
4.1. Filtering Data from A High-Resource Language for Translation and Test Generation
Filtering Step #Functions
All functions 22,311,478
With docstrings 5,359,051
Typechecked and returns value 459,280
No TODOs and no benchmark solution 432,361
Test generation 157,767
90% line coverage from tests 133,168
Table 1. Size of the Python source dataset after each filtering step.
The first step in MultiPL-T is to filter code from a high-resource language to serve as the translation source for our semi-synthetic data. We use Python because it has the highest representation in The Stack and because MultiPL-E can compile Python function signatures and test cases to several low-resource languages. However, our approach could easily be adapted to work with other high-resource languages.
Filtering Python Functions Before Translation
The Stack has 22 million Python functions (Table 1). However, not all of these are amenable to translation and test-based validation with MultiPL-T. One could naively try to translate and generate tests for all 22M functions. However, since doing so requires GPUs, it would be prohibitively expensive. Instead, we aggressively filter the 22M functions down to ~400,000 functions using the following steps:
  1. (1)
    We exclude Python functions that do not have a docstring or use non-ASCII characters. One could generalize to include functions that have an associated comment. However, we still end up with up over 5M candidate functions with this simple filter.
  2. (2)
    We use the Pyright (Pyright, 2023) Python checker to validate that each function returns a value, uses only the Python standard library, and is thus likely type-correct [^6] Pyright is a heuristic type-checker that does not guarantee soundness.. This narrows the 5M functions to approximately 460,000 functions.
  3. (3)
    We exclude Python functions that have comments suggesting the implementation is incomplete (e.g. “TODO”). It turns out that a fair amount of code on The Stack is incomplete; these functions are not likely to be useful training data. To avoid data contamination, we filter out functions whose prompt or solution appears in widely-used Code LLM benchmarks by finding exact matches of the prompts (Chen et al., 2021; Austin et al., 2021).
The final dataset contains 432,361 Python functions. With this narrower set of functions, we move on to the next steps that require GPUs.
(a) CDF of test suite sizes
(b) CDF of test suite coverage
Figure 6. Python test suite sizes and coverage distribution among the unfiltered test generation dataset
def count(word):
    """Count the number of Xs in a word."""
    count = 0
    for letter in word:
        if letter == "Y":
            count += 1
   return count
Figure 7. An example prompt to generate a unit test where the code and comment are inconsistent: the comment counts X’s, but the code counts Y’s. Using a Code LLM to generate tests helps expose the inconsistency: we get some tests that are based on the comment and others based on the code.
4.2. Generating Python Unit Tests
The next step in MultiPL-T is to generate unit tests for each Python function. We will then compile these unit tests to target low-resource languages using the test case translators from MultiPL-E. However, instead of using a traditional test generator that synthesizes tests from code (as in (Lukasczyk et al., 2020)), we use a Code LLM to generate test cases by prompting the model to produce an assertion. The Code LLM conditions on the source code text and thus serves as a weak detector for inconsistencies between code and comments. For example, figure 7 shows a function that would be a bad training item: the code counts Y’s, but has a comment that says it counts X’s. When we generate several test cases independently, we end up with tests for both X and Y, which lowers the ratio of passing tests. Conditioning on the text makes it less likely that inconsistent tests will be generated, decreasing the likelihood that functions will be filtered out of the training set due to low coverage (as described below).
We prompt StarCoder-15B to generate five independent test suites for each function at high temperature [^7] Temperature is a parameter that controls the randomness of the next-token predictions. A higher temperature results in more varied (less predictable) output, while a lower temperature produces more conservative (more predictable) results. (0.8) to get a diverse set of candidate tests. We parse each generated test suite and extract all test cases that are suitable for translation using MultiPL-E. We take the set of matching test cases and run each test in isolation in a container to verify that it passes, discarding any that fail. If no correct tests are generated, we discard the function. The result is a set of nearly 160,000 Python functions with at least one passing test case. Figure 5(a) shows the distribution of test suite sizes. The median number of test cases per function is 7, and the mean is 12.1.
Filtering on test coverage
Given the dataset of Python functions with docstrings and test suites, our next step is to filter out functions with low test coverage. We use line coverage as the coverage metric, and exclude all functions with less than 90% line coverage. The result is a dataset of 133,668 functions with 90% line coverage from tests.
Since we start with nearly 160,000 functions, this implies that most of the generated test suites that work have high line coverage. In fact, most functions have 100% line coverage (Figure 5(b)). This stringent criterion ensures that the functions in our final set are not just correct but are also comprehensively tested, reinforcing the reliability of our dataset.
Type inference
The steps described above are sufficient to generate data for an untyped low-resource language (e.g., Racket or Lua). However, for a typed target (e.g., OCaml or Julia), we also need to infer types. [^8] Julia can be used with or without types, and our dataset has both typed and untyped examples. Even when the source language supports type inference, we need to infer Python types to translate test cases. For example, a Python function that consumes an optional integer can receive None or a number $n$. However, when we translate that function to OCaml, we must translate that number to Some($n$). We can only do this by inferring that the Python type is Optional[Int].
Our approach to type inference is simple: we deduce types based on test cases, ignoring the function body. We extract the instance type of each argument and expected return value in each test, computing the union type between the types at the same position among tests. For example, if the test cases apply foo(1) and foo(None), we infer Union[int, None] as the type of foo’s argument. Moreover, we simplify Union[T, None] to the more canonical Optional[T]. In the example above, Union[int, int, None] would be then simplified to Optional[int].
Following the steps above produces two datasets of Python functions—one with and one without type annotations—where every function has a docstring, and a suite of unit tests that achieve high coverage. These datasets can be used to generate training data for any low-resource language.
4.3. Translation from a High-Resource to a Low-Resource Language
Given a dataset of commented Python functions with high coverage unit test suites, our next goal is to translate the dataset from Python to a target language $L$ and use tests to validate the translation.
Translation with a Code LLM and MultiPL-E
We use a modified version of MultiPL-E (Cassano et al., 2023) to translate each Python function into an equivalent function in the target low-resource language. We construct a MultiPL-E prompt with the following three parts:
  1. (1)
    Docstring: We turn the Python docstring into a comment in the target language. The MultiPL-E toolchain translates between different comment formats, and also alters common type names in natural language using simple rules. For example, when translating from Python to OCaml, we turn “dictionary” into “association list”.
  2. (2)
    Function signature: We turn the Python function signature into a function signature in the target language. This involves translating types from Python into the target language and is trivial for untyped targets.
  3. (3)
    Original Python code: Finally, we add a comment (in the target language) that contains the original Python code. We find that this additional information increases the chance that the model generates a correct translation (§ 5.5).
Figure 5 highlights an example prompt and test suite for translating a descending sort function written in Python to OCaml in the programs labeled 1 and 2. MultiPL-E translates comments written in Python to OCaml and translates each test case and the function signature from Python to OCaml. The original Python code is added as part of the comment.
Given this prompt, we use StarCoderBase-15B to generate translations of each problem in our Python dataset. For all of our languages, we generate 50 translations [^9] For OCaml, we generated 100 translations per problem, as the base pass rate was significantly lower than other languages. at high temperature (0.8), to encourage the Code LLM to produce a more diverse set of candidate solutions (Chen et al., 2021).
Checking translations with compiled tests
A Code LLM is quite likely to produce faulty translations; in our case, this is even more likely, since we specifically target languages on which the Code LLM performs poorly. We address this problem by translating test cases from Python to the target language using a simple, recursive compiler. MultiPL-E has a suite of compilers that translate simple Python assertions into assertions in 20+ other programming languages. The compilers support assertions that are simple equalities between first-order values, specifically atomic Python data and collections (lists, tuples, and dictionaries). We use these compilers to translate tests from Python to each target language, removing test cases that MultiPL-E does not support. If we are left with zero test cases, we discard the function entirely.
Given the set of 50 generated translations for each function, we select only those solutions that pass all tests. This may include selecting several solutions to the same problem, which is beneficial to the model in terms of learning diverse code styles.
4.4. Deduplication
We define a set of solutions to the same problem as diverse when they are distinct in form but behave the same on the test suite. Ensuring diversity in our generated and verified solutions is critical for ensuring that fine-tuned models learn a range of syntactic and semantic features in the target programming language. Moreover, redundant or similar solutions can dilute a dataset’s efficacy (Lee et al., 2022).
Algorithm 1 Parallelized ROUGE-L deduplication procedure.
[…]
2:     $G\leftarrow$ GroupByPrompt($I$) $\triangleright$ Group items by prompt (comment)
[…]
4:     for $i\leftarrow 0$ to $rounds$ do
[…]
7:     end for
[…]
9:end procedure
10:procedure DedupGroups($G$, $t$) $\triangleright$ Deduplicates item within each group in the list of groups ($G$).
11:     $\ell\leftarrow[]$
12:     for $g\leftarrow G$ in parallel do
[…]
14:         for $i\leftarrow 0$ to $\mid g\mid-1$ do
15:              for $j\leftarrow i+1$ to $\mid g\mid$ do
[…]
17:                       continue
18:                  end if
19:                  $a,b\leftarrow$RemComments($g[i]$), RemComments($g[j]$)
[…]
22:                  end if
23:              end for
24:         end for
[…]
26:     end for
27:     return $\ell$triangleright A list of items (ungrouped)
28:end procedure
Just resampling with high temperature does not guarantee diverse solutions: the LLM may still produce solutions that are nearly identical (e.g., with a few variables renamed). To address this, we employ a deduplication algorithm based on ROUGE-L (Lin and Och, 2004). ROUGE-L is a metric of text summarization quality, and quantifies the syntactic overlap between two pieces of text with a score ranging between 00 and $1$ where $1$ indicates highest similarity. We use a 0.6 as the similarity threshold for discarding duplicates. Before comparing a pair of solutions, we remove all comments from the code, as it may introduce noise in the deduplication process.
Running ROUGE-L on all pairs of items is prohibitively expensive. Instead, we use a heuristic that is amenable to parallelization (Algorithm 1). We apply ROUGE-L to deduplication items in fixed-size groups (we use size 200). Initially, the groups are solutions to the same prompt. We then randomly regroup items and run grouped deduplication again. The number of rounds of deduplication is proportional to the total number of items: more rounds increases the likelihood that duplicates will be removed. Ultimately, this results in a set of diverse, accurate, and semantically equivalent solutions for each prompt.
5. Evaluation
We use MultiPL-T data to fine-tune a large, medium, and small-sized Code LLM: CodeLlama-34B, StarCoderBase-15B, and StarCoderBase-1B. Both StarCoder models have the same architecture and were trained on exactly the same data. However, StarCoderBase-1B is significantly cheaper to train and evaluate, so we use it for the majority of the results in this section. In contrast, CodeLlama-34B is a larger model that was trained on an undisclosed dataset of publicly available code. We use it to demonstrate that our results generalize to larger Code LLMs.
Training hyperparameters
We fine-tune all models with a sequence length of 2,048 tokens. StarCoderBase-1B is fine-tuned for seven epochs with batch size $8$, with learning rate $3\times 10^{-5}$, $10$ steps warmup, and cosine learning rate decay. For StarCoderBase-15B and CodeLlama-34B, we make these configuration changes: ten epochs, batch size $32$, and learning rate $2\times 10^{-5}$.
Estimated computing resources used
The work for this article was done over several months using V100 (32GB), A100 (80GB), and H100 (80GB) NVIDIA GPUs, as they were available across several clusters and servers. We estimate that we spent nearly 500 days of A100 (80GB) GPU time with the following breakdown:
  • Training: ~1,624 hours fine-tuning several versions of Code Llama-34B, StarCoderBase-15B, and StarCoderBase-1B. These include the models presenting in this section and in § 3.
  • Evaluation: ~230 hours running benchmarks, which include MultiPL-E and the new adversarial benchmark (§ 5.3).
  • Dataset Generation: ~9,984 hours generating the MultiPL-T training sets. This was the most significant use of resources, but is reusable for future model development.
Training and evaluation tools
We experimented with several technologies during development. The final MultiPL-T pipeline uses vLLM (Kwon et al., 2023) for inference, DeepSpeed ZeRO to fine-tune larger models (Rajbhandari et al., 2020), Transformers (Wolf et al., 2020), and MultiPL-E (Cassano et al., 2023) with several modifications, such as supporting OCaml.
Language StarCoderBase-1B StarCoderBase-15B CodeLlama-34B
Base Fine-tuned Base Fine-tuned Base Fine-tuned
OCaml 1.5 9.7 6.9 19.9 18.3 27.4
Racket 4.7 11.3 11.8 21.0 15.9 29.1
R 5.4 8.9 10.2 17.3 18.2 25.5
Julia 11.3 15.6 21.1 35.2 31.8 43.5
Lua 12.1 17.3 26.6 31.0 38.1 43.9
Table 2. MultiPL-E pass@1 scores for 1B, 15B, and 34B parameter models before and after fine-tuned on MultiPL-T data. The fine-tuned models show significant improvement.
5.1. Fine-Tuning Small, Medium, and Large Code LLMs with MultiPL-T
Fine-tuning our base models with MultiPL-T-generated data on our target languages improves performance on MultiPL-E across the board (Table 2). For each model, we fine-tune a separate model for Julia, Lua, OCaml, R, and Racket. We checkpoint and evaluate the models at each epoch and report the peak performance. It is not a goal of this paper to maximize MultiPL-E scores. In fact, the next section suggests that it would be easy to improve some of these scores by either training longer on the data we already have or by letting MultiPL-T generate more data. Beyond the generally improved performance, we draw several other conclusions from our evaluation (Table 2):
  1. (1)
    For Racket and OCaml, which are the lowest-resource languages that we evaluate, the fine-tuned 1B parameter models perform as well as StarCoderBase-15B at baseline.
  2. (2)
    Racket and OCaml show the largest relative gains. For example, the fine-tuned versions of both models have more than double the score of their base models, with particularly large gains for OCaml.
  3. (3)
    Lua obtains relative gains of 42% for 1B and 17% for 15B. However, these gains are significant and put the fine-tuned models’ Lua performance on par with the base models’ performance on the highest-resource languages. For example, StarCoderBase-15B achieves $30.6$ on MultiPL-Python, and our fine-tuned model achieves $31.0$ on MultiPL-Lua.
  4. (4)
    Julia also shows a significant relative gain of 67% for 15B, achieving a score on MultiPL-Lua that exceeds the base model’s MultiPL-Python scores.
  5. (5)
    We also fine-tune and evaluate CodeLlama-34B and see significant improvements as well.
Overall, when compared to open models that are not fine-tuned on proprietary data, these fine-tuned models achieve state-of-the-art benchmark scores on Julia, Lua, OCaml, R, and Racket at 1B, 15B, and 34B model sizes.
Figure 8. We fine-tune three versions of StarCoderBase-1B on 25k MultiPL-T generated training items. The $y$-axis measures performance on MultiPL-E and the $x$-axis counts the number of tokens. The points on the line mark epochs. Racket is far less verbose than Lua or OCaml, and thus has fewer tokens at each epoch. Epoch 0 is the base model.
5.2. Fine-Tuning Efficiency
We now investigate how StarCoderBase-1B’s performance on MultiPL-E varies during fine-tuning. On three languages (Lua, OCaml, and Racket), we fine-tune StarCoderBase-1B for seven epochs of MultiPL-T data and evaluate at each epoch. However, the datasets that we have generated are imbalanced in the number of training items (Table 3). To better balance the training sets, we randomly sample 25,000 items from each dataset to get similarly-sized fine-tuning sets for each language. [^10] However, small differences remain: Lua is more verbose than OCaml, so 25,000 training items for Lua is more data than 25,$0$ training items for OCaml.
Language Size
R 37,592
Racket 40,510
OCaml 43,401
Julia 45,000
Lua 48,194
Table 3. Dataset sizes.
In figure 8, we see that performance increases substantially after a single epoch of MultiPL-T data for the lowest resource languages (Racket and OCaml). However, a higher-resource language (Lua) requires more data to realize even modest gains. These results are what one would expect: lower-resource languages enjoy easy and efficient gains from fine-tuning with MultiPL-T data.
On the other hand, MultiPL-T requires more computing resources to generate data for low-resource languages, so the overall efficiency gap between languages is narrower than the figure suggests: it does not show the cost of generating training data. Nevertheless, MultiPL-T data only needs to be generated once for a given language and can then be reused to train many models (§ 5.1).
Language StarCoderBase-1B StarCoderBase-15B
Base Fine-tuned Base Fine-tuned
OCaml 33.0 33.7 50.6 42.9
Racket 19.3 22.7 28.4 41.3
Lua 26.3 46.9 48.7 51.3
Table 4. The pass@1 scores of the original models and fine-tuned models on our new adversarial benchmark that uses user-defined types, higher-order functions, helper functions, and non-standard libraries. The scores suggest that the 15B model may have overfit when fine-tuned on OCaml. However, the other models do the same or better after fine-tuning.
5.3. A New Adversarial Benchmark
A limitation of the MultiPL-T datasets is that every training item is a single function without any other context: they may use standard libraries, but cannot depend on other functions, classes, or third-party libraries. Thus it is possible that fine-tuning a model on MultiPL-T data will make it overfit to this format. [^11] Recall that the base models have been pretrained on natural code that is not constrained to the MultiPL-T format. Thus appropriate fine-tuning should not eliminate their ability to generate code that is not in the MultiPL-T format. Moreover, conventional Code LLM benchmarks, including MultiPL-E, will not expose this problem, because the benchmark tasks largely involve generating standalone functions that use only standard libraries.
To determine if this kind of overfitting is a problem, we construct a new, multi-language benchmark with fourteen problems. Every problem has hidden test cases and a prompt that exercises the model’s ability to use user-defined types, higher-order functions, helper functions, or external libraries (Table 5). We manually translate these prompts to idiomatic OCaml, Racket, and Lua.
We evaluate StarCoderBase 1B and 15B on this new benchmark (Table 4). The results suggest that the 15B OCaml-tuned model may have overfit to the MultiPL-T format. However, all other fine-tuned models do the same or better. We speculate that fine-tuning on a mix of natural and MultiPL-T data will decrease the likelihood of overfitting.
Problem Description
Add Left of NumTree
Context: a type for a tree of numbers. Problem: add the numbers on the left branches.
Add Subway Station
Context: a subway system represented as a graph. Problem: add connections between stations.
Map Over Android Phones
Context: a type for phone models. Problem: a mapping function that only applies to Android phones.
Electrify Instruments
Context: a type hierarchy for several musical instruments with flag that indicates if they are electric. Problem: mark instruments as electric.
Collatz Depth
Context: empty. Problem: requires a solution to see how far a number is from converging on 1 in the Collatz sequence.
Mirror a Tree
Context: a type for trees. Problem: tree reflection.
Double Do It
Context: 2 of 3 helper functions. Problem: define third helper and primary function.
Points and Lines
Context: types for point and lines types Problem: Manhattan distance.
Series
Context: a type which represents a series with its current value and update function. Problem: define a function which updates to the next value in the series.
Shapes
Context: several shape types and functions/methods. Problem: define a function that uses the helpers.
Social Time
Context: types that define calendar events. Problem: calculate time spent on a particular kind of event.
Decode Message
Context: imports a common cryptography library. Problem: Decode AES encrypted message.
Verify Source
Context: imports a common cryptography library. Problem: verify signature.
Start AES
Context: imports a common cryptography library. Problem: generate an AES key and encrypts it.
Table 5. The 14 adversarial problems.
5.4. Evaluating Coding Style
Category Grading Item Max Deduction
Text Dangling parentheses $-0.5$
Line too long $-1$
Using car/cdr $-0.5$
cond with round brackets $-0.5$
Definitions let expression not at beginning of function body $-1$
Nesting define or let expressions $-2$
Unecessary use of let* or letrec $-1$
Defining useless local variables $-1$
Not defining helpers or variables for reused code $-1$
Conditionals Nested if expression instead of cond $-2$
Using (if COND #t #f) $-1$
Traversal Using iteration when recursion is available $-3$
Figure 9. Our Racket style rubric for grading generated programs. The maximum score for a program is 15 points. Items are grouped into categories according to the type of error they entail.
Figure 10. Results of our Racket style evaluation for 30 out of 35 selected HumanEval problems. Omitted problems have no deductions and both models scored 15/15 points. Deductions for each item category are averaged across graders for each problem. Fine-tuning has a small negative impact on coding style.
A potential limitation of MultiPL-T is that it may negatively impact the style of generated code, since the training items are translated from Python. We study this issue in Racket using a qualitative evaluation process. We developed a Racket style grading rubric (Figure 9) based on the Racket style guide and our experience teaching and grading Racket programming assignments. The rubric outlines grading items and their corresponding deductions. Several deductions are designed for style problems that arise in code written by beginning students, such as needlessly long lines. Others penalize Lisp style, such as using car and cdr instead of first and rest. Finally, the largest deduction is for using imperative iteration when a simple functional solution is possible.
We use our rubric to grade the HumanEval solutions produced by StarCoderBase-1B before and after fine-tuning on MultiPL-T Racket data. We grade only the 35 problems that both models are able to solve at least once (thus we omit several problems solvable only after fine-tuning). Since we generate several solutions for each problem, we select the most common working solution, and in case of ties, we select the one with the best style. We grade 70 Racket programs in total: 35 produced by the base model and 35 after fine-tuning. To avoid bias, we use two graders and anonymized the selection and grading process by assigning random IDs to each program and tracking their provenance on a hidden spreadsheet. The two graders have substantial Racket teaching experience and we find minimal discrepancy between them: their scores differed by more than 1 point for just 15 out of the 70 candidate programs, and only differed by 3–4 points for two programs.
We compute the mean overall score for the base model and MultiPL-T model over the 35 HumanEval problems. We find that the base model achieves a style score of $89.5\%$ and the fine-tuned model $85.2\%$. In other words, the mean grade for a base model program is $13.4/15$ while for a fine-tuned model it is $12.8/15$. Thus fine-tuning leads to a slight decrease in our Racket style score.
We inspect the 17 programs that scored higher for the base model than the fine-tuned model. We find that the fine-tuned model is more likely to use nested if expressions in these programs, as well as performing iteration where recursion is available. Conversely, we inspect the eight programs that scored higher in the MultiPL-T model and found that the base model is more likely to use direct recursion with car/cdr while the fine-tuned model uses Racket list abstractions. figure 10 shows the breakdown of the kinds of deductions assigned per problem to each model, where deductions are averaged among graders.
The results of our evaluation are consistent with how we generated our training data from Python code. Overall, although fine-tuning slightly decreases the model’s ability to generate idiomatic Racket code, it also significantly increases its ability to generate correct Racket code as we showed earlier. Moreover, both models perform well on the Racket style rubric, which suggest that the tradeoff is minimal.
Language Basic With Canonical
OCaml 26.1 23.9
Racket 34.7 56.8
Lua 51.4 68.5
Table 6. The pass@50 performance of the basic prompt compared to the prompt with the original Python code for OCaml, Racket, and Lua.
5.5. Translation Prompt Ablations
The final part of our evaluation concerns the efficiency of the MultiPL-T pipeline, and not the quality of LLM-generated code. MultiPL-T uses StarCoder-15B to translate training items from Python to low-resource languages. The success rate of this translation is dependent on both the quality of the pretrained model (which is fixed) and the quality of the prompt (which we design). We tried several prompt variations during development, and eventually settled on a prompt that includes the original Python code in a comment (§ 4.3).
Doing a complete ablation with all five languages and ~133,000 functions would be prohibitively expensive. Instead, we do an experiment with a random sample of 1,000 source Python functions. We use the LLM to translate these 1,000 functions to OCaml, Racket, and Lua with two different prompt formats: with and without the original Python source. We generate 50 candidates for each prompt. [^12] Thus this small experiment still requires 300,000 generations from the LLM and takes about 1/2 a day on an A100 GPU. We evaluate performance using pass@50, which is the likelihood that the model produces at least one correct solution in 50 attempts.
As shown in table 6, adding the original Python to the prompt substantially increases the likelihood of a successful translation to Racket and Lua, but slightly decreases the likelihood of a successful translation to OCaml. We can only speculate about why this happens: Python may be misleading the model and OCaml seems further removed from Python than Racket or Lua.
6. Discussion
We have shown that MultiPL-T is an effective and efficient method for generating semi-synthetic training data for low-resource programming languages. In this section we discuss the implications of extending MultiPL-T in various ways.
Generalizing to other programming languages
We hope it is clear to the reader that the MultiPL-T approach is straightforward to generalize to more programming languages. The language-specific work involves 1) translating comments and function signatures into an appropriate prompt, and 2) writing a compiler that can translate simple assertions from the source to target. MultiPL-E already supports both these steps for 20+ programming languages, several of which are low-resource, including D, Bash, MATLAB, Haskell, and Perl. So, generating fine-tuning sets for these languages may just be a matter of running the MultiPL-T pipeline for a few days on GPUs.
Generalizing to other LLMs
Although this paper fine-tunes the Code Llama and StarCoder family of Code LLMs, our datasets could also be used to fine-tune other LLMs. We use StarCoderBase-15B to generate data, but it would be straightforward to swap in some other model as well.
“No-resource” languages
MultiPL-T targets low-resource languages, but it is unlikely to work as-is for “no-resource” languages that StarCoderBase is not trained on at all. It may be possible to cleverly prompt the model with enough information about a no-resource language to bootstrap data generation. But, doing so efficiently may be challenging.
Composability with self-instruct
Although we have argued that self-instruct is unlikely to succeed on a low-resource language, self-instruct and MultiPL-T could be composed together in a natural way: one could generate a high-quality dataset of instructions in a high-resource language, and then use MultiPL-T to translate them to a low-resource language. Given the effectiveness of WizardCoder (Luo et al., 2023) at Python, this composition seems likely to succeed.
7. Related Work
Code translation with language models and unit tests
A number of projects use language models to translate code between programming languages and test that the generated translations are correct by compiling working unit tests from one language to another. TransCoder-ST (Roziere et al., 2021) and CMTrans (Xie et al., 2024) uses these techniques to generate training data for a code translation model between Java, Python, and C++, whereas MultiPL-E (Cassano et al., 2023), MBXP (Athiwaratkun et al., 2022), and BabelCode (Orlanski et al., 2023) translate Code LLM benchmarks from Python to 10+ programming languages. A distinguishing feature of MultiPL-T is that it employs an off-the-shelf pretrained Code LLM (StarCoder) to both generate test cases and translate code to low-resource languages. When the aforementioned papers were written, the best open Code LLMs were far less capable than StarCoder: they were trained on fewer programming languages using far less training data, and they were an order of magnitude smaller. Thus we believe the MultiPL-T approach would have likely failed. Although capable closed models were available, they were either rate limited or prohibitively expensive for the scale of data generation that MultiPL-T needs. For example, Cassano et al. (2023) report that they used a commercial model during a free beta period, but it would have cost \$37,000 with equivalent released models. MultiPL-T’s data generation requires an order of magnitude more queries, making it prohibitively expensive to use with commercial models at 2023 prices.
Instruction tuning
To get an LLM to perform a desired task, the user must prompt it in the right way. There are several techniques for instruction tuning LLMs to better follow natural, human-written instructions. One approach uses human annotators to write sample instructions and give feedback on a large number of model generations (Ouyang et al., 2022), but this is expensive and requires significant resources. A cheaper approach is to have a capable LLM self-instruct to generate instructions from a relatively small set of human-written seed instructions (Wang et al., 2023). Evol-Instruct uses an LLM to create variations of instructions (Luo et al., 2023). These techniques have been used to create datasets for instruction-tuning Code LLMs (Chaudhary, 2023; Muennighoff et al., 2023; Luo et al., 2023). These datasets focus on high-resource languages, and, as we show in § 3.2, they are unlikely to succeed for low-resource languages.
Training on high quality data
Training on high-quality data is an effective way to reduce both the size of an LLM and the volume of training data needed, while maintaining performance. Gunasekar et al. (2023) achieve high HumanEval scores on a small model with a modest amount of “textbook quality” training data. This includes both natural and synthetic data generated by a more capable model. Their work targets Python, and we argue in § 3 that the approach is less likely to succeed on low-resource languages.
Proprietary Code LLMs
At the time of writing, there are proprietary LLMs that perform better at programming tasks than the open models we build upon (OpenAI, 2023a; Anthropic, 2023a; Anil et al., 2023). However, most of these models only support inference (i.e., running the trained model) and not training or fine-tuning. Even when fine-tuning is possible, because these models are trained on closed data sets, we would not be able to compare MultiPL-T to the natural baseline of training longer on existing data. Moreover, a significant limitation arises from the proprietary licensing constraints of these models. Many of their licenses expressly forbid the use of generated data to train other models (OpenAI, 2023b; Google, 2023; Anthropic, 2023b).
Models fine-tuned on the output of proprietary Code LLMs
Despite the fact that many proprietary models forbid using their generated data to train other models, there are models that do so. The Phind and WizardCoder (Luo et al., 2023) projects have fine-tuned both StarCoder and Code Llama models on data generated by GPT-3.5 and GPT-4. Phind also uses additional proprietary data that is orders of magnitude larger than the MultiPL-T training data. These models outperform our fine-tuned Code Llama on R, and are close on the other programming languages.
However, we do not believe these comparisons are very meaningful. One could fine-tune these models further on the MultiPL-T data we already have, or even use them (or GPT-4) to translate Python to low-resource languages. But, we would not be able to compare to the natural baseline of training longer on existing data, which for these models is the proprietary OpenAI training set. The strength of the MultiPL-T approach is not that it does better in an absolute sense, but that fine-tuning with MultiPL-T data is better than training longer on existing data. We show that this holds for StarCoderBase and its training set, and can only do so because both are open.
8. Conclusion
In the last few years, Code LLMs have rapidly made their way into more and more programming tools. However, the quality and reliability of Code LLMs is highly language-dependent: they can be remarkable on high-resource programming languages, but are far less impressive at working with low-resource languages. It is possible that in the near future, a large number of programmers will expect LLM-based technology to work, just as many programmers today expect syntax highlighting, continuous analysis, or type-based completion in their programming environments. We hope that MultiPL-T—a methodology for generating large-scale, high-quality fine-tuning datasets in low-resource languages—will help low-resource languages compete in a world where many developer tools rely on Code LLMs.
The MultiPL-T fine-tuning data (and code) are also open: they are constructed from the StarCoder training data (The Stack) and augmented by StarCoder itself. We deliberately do not use a more capable proprietary model to fine-tune an open model. This allows us to demonstrate that fine-tuning on MultiPL-T data is more effective and efficient than training longer on existing data. We evaluate MultiPL-T in several other ways, including a new adversarial benchmark and a qualitative evaluation of coding style. When compared to open models that are not fine-tuned on proprietary data, MultiPL-T achieves state-of-the-art results for Julia, Lua, OCaml, R, and Racket.
Appendix A Human-Written Prompt for Self Instruct
We use the following five functions to start a round of self-instruct for Racket. We generate results at temperature 0.8.
;; watching-you?: Number -> Boolean
;; Sees if the number has a "00" in it
(define (watching-you? num)
    (string-contains? (number->string num) "00"))
;; add-odds: [List-of Numbers] -> Number
;; Adds all the odd numbers in a list
(define (add-odds lon)
    (foldr + 0 (filter odd? lon)))
;; repeat: String Number -> [List-of Strings]
;; Repeats a string num amount of times
(define (repeat str num)
    (if (= num 0) ""
    (string-append str (repeat str (- num 1)))))
;; sum-remainders: [List-of Numbers] Number -> Number
;; Sums the remainder of the numbers in the list when divided by num
(define (sum-remainders lon num)
    (foldr + 0 (map (lambda (n) (remainder n num)) lon)))
;; is-palindrome?: Number -> Boolean
;; Checks if a number is a palindrome
(define (is-palindrome? num)
    (equal? (number->string num)
            (list->string
                (reverse (string->list (number->string num))))))
Appendix B More Details on Evaluation
Language StarCoderBase-1B StarCoderBase-15B CodeLlama-34B
Epoch
OCaml 3 4 2
Racket 6 8 4
Lua 6 4 2
R 6 2 4
Julia 4 6 4
Table 7. Epochs chosen for each language and model fine-tuned as described in § 5.
The results in table 2 and table 4 report the best performance of the fine-tuned models on MultiPL-E, which we find at the epochs highlighted in Table 7.
B.1. Evaluation On MultiPL-MBPP
Language StarCoderBase-1B StarCoderBase-15B CodeLlama-34B
Base Fine-tuned Base Fine-tuned Base Fine-tuned
OCaml 3.7 20.2 17.2 28.9 28.9 35.8
Racket 4.7 19.3 21.1 30.5 29.8 39.7
Lua 18.7 26.6 35.2 40.3 42.8 47.4
Table 8. MultiPL-MBPP pass@1 scores for 1B, 15B, and 34B parameter models fine-tuned on MultiPL-T data.
The main body of this article focuses on the MultiPL-E translation of HumanEval, but MultiPL-E also translates MBPP (Austin et al., 2021), which is a program synthesis benchmark comprised of 427 crowdsourced Python problems. This benchmark is similar to MultiPL-E, but more than twice the number of problems and is typically considered to be less challenging. We further evaluate our fine-tuned models for OCaml, Racket, and Lua MultiPL-MBPP.
Appendix C Pass Rate with Adversarial Benchmark
Comparison of StarCoderBase-15B models’ performance when they were just trained on the Stack versus after being fine-tuned on MultiPl-T dataset.